### v3

@@ Abstract
Large Language Models (LLMs) are increasingly deployed as interacting agents, forming “LLM societies”. Understanding whether these societies can self-organize and coordinate on norms without external influence is crucial to understand their risks and opportunities. Here we explore their opinion dynamics finding that it is governed by a majority force coefficient such that LLM societies can spontaneously reach consensus only up to a critical group size. This critical size grows exponentially with the language understanding capabilities of the models, exceeding the typical size of informal human groups for advanced LLMs. These results reveal emerging self-organization properties in LLM societies and provide insights for designing collaborative AI systems where coordination is either a goal or a risk.
Introduction
@@
Large Language Models (LLMs) have proven individual capabilities for a wide range of applications, such as summarization [1], sentiment analysis [2, 3], scientific research [4] or mathematical reasoning [5]. Agents driven by LLMs can be used in group settings where several agents interact with each other in collaborative tasks [6, 7, 8]. Collaboration setups where multiple LLMs have different roles and tasks, such as AutoGPT  [^1] https://github.com/Significant-Gravitas/AutoGPT and more recently Microsoft’s AutoGen [9] or OpenAI SWARM [^2] https://github.com/openai/swarm, are qualitatively different from ensembling techniques, where interaction between different models is absent [10]. Recent advancement on on-device LLMs are also leading to AI-powered devices and assistants, such as Siri [^3] https://openai.com/index/openai-and-apple-announce-partnership/ or the Humane AI Pin, that can perform everyday tasks in interaction with each other, for example coordinating events or negotiating prices.
@@
As we move towards a society where LLM agents interact with each other in our behalf and can coordinate, it becomes important to understand their ability to agree with each other in large groups. This can motivate new applications but also help identifying risks stemming from undesired collective behavior. For example, trading bots interacting through the stock market can lead to flash crashes [11]. Current research on the behavior of LLMs has mostly focused on their behavior in isolation [12, 13, 14, 15, 16, 17] and collective behavior has been explored less [18, 19, 20], mostly with a focus on social simulation of network structures [21, 22, 23], opinion and information spreading [24, 25] and online interaction [26, 27]. To be prepared for large numbers of interacting LLM agents, we need to understand if they can display emerging consensus, what determines the abilities of LLM agents to coordinate, and at what scale this can happen.
@@
Group coherence is related to coordination, as the agreement on shared norms creates a social contract that regulates behavior [28]. Norms are also a necessity even in situations where there is no information about the quality or utility of any option, for example in the case of animal coordination when moving collectively [29]. Across species, this leads to a scaling of the average group size with brain structure, with human groups reaching sizes between 150 and 300 [30, 31], as documented by archaeological records [32] and contemporary experience [33]. To reach larger scales, human societies have built institutions and other ways of decision making, but the cognitive limit of about 250 contacts remains even in an online society [34, 35]. In this paradigm, language can work as a tool to learn about the behavior of others in a more efficient way than through observation, for example through gossiping [36, 37], thus leading to the hypothesis that language abilities are a factor in the capability of LLM societies to reach consensus. These are the new questions of AI anthropology, where the insights and methods for the study of animal and human societies can be applied to study the complexity of LLM societies.
@@
In this article, we investigate the ability of LLM societies to reach consensus about norms for which there is no information supporting one option over another. The emergence of consensus is a foundational aspect of social systems, where individual interactions lead to the formation of a unified agreement or shared understanding without the need for a central authority or structure [38, 39]. We develop a framework to test if LLM societies can reach consensus and use it to analyze a benchmark of proprietary and open-source models. We apply insights from previous opinion dynamics research to measure a majority force that determines the possibility of consensus in LLM societies. This majority force is a function of language understanding capabilities of models and of group size, with consensus not emerging beyond a critical size for a given LLM. Furthermore, we find that some of the most capable LLMs are able to reach consensus at scales beyond what human groups typically achieve.
@@ Results
Opinion Dynamics of Large Language Models
To investigate the coordination abilities of LLM societies, we perform simulations using agents guided by various LLMs, belonging to the GPT, Claude, and Llama families. Simulations run as follows (see Methods for more details). Each agent is assigned an initial opinion randomly chosen from a binary set (e.g., ”Opinion A” and ”Opinion B”), where the first is chosen as the norm for reference. At each time step, a single agent is randomly selected to update their opinion. The selected agent receives the list of all other agents with their current opinions and is then prompted to choose their new opinion based on this information. This approach mirrors binary opinion dynamics such as the voter model or Glauber dynamics [40], where agents update their opinions based only on peer interactions. However, unlike traditional Agent-Based Models with predefined opinion update rules and equations, here agents autonomously decide their opinions based on their LLM.
To characterize the evolution of the system, we define the average group opinion
@@
Here $s_{i}$ is the opinion of agent $i$, the first opinion is $s_{i}=+1$ and the second is $s_{i}=-1$ (i.e., in favor and against the norm). $N_{+}$ and $N_{-}$ are the number of agents supporting the first and second opinion respectively, while $N$ is the total number of agents. In these terms we can define the consensus level $C=|m|$ that quantifies the level of agreement among agents. Full consensus corresponds to $C=1$, while $C=0$ means that the system is split in two groups of equal size and opposite opinion. Three scenarios can occur in the evolution of $C(t)$:
  • $C$ can converge to 1 ($m$ converges to ±$\pm 1$), meaning that all agents coordinate and consensus is reached.
  • $C$ can fluctuate around a value greater than 00 without ever reaching $1$. In this case only a partial consensus is reached for a subset of the group.
  • If $C$ keeps fluctuating close to zero, consensus is completely absent and the group is constantly in a disordered state.
Figure 1: Opinion Dynamic of LLM agents. Left: Evolution of the consensus level over time for five different models and group size $N=50$. The box plots on the right show the final consensus level over $20$ simulations. Some models always reach consensus, while the others never do so. Right: adoption probability $P(m)$ as function of the collective opinion $m$ in the group. Solid lines are fits of the curve $P(m)=0.5[\tanh(\beta\cdot m)+1]$ to the empirical data. The inset shows rescaled probabilities $P(\tilde{m})$ ($\tilde{m}=\beta\cdot m$) and confirms that all LLMs follow the same universal function.
We show in Fig. 1 the evolution of the consensus level in societies of $N=50$ LLM agents and five different models, where the boxplot shows $C(t=10)$ over $20$ realizations. The two most advanced models considered, Claude 3 Opus and GPT-4 Turbo, reach consensus in all simulations, which corresponds to $|m|=1$ in the boxplot. On the other hand, less advanced models (Claude 3 Haiku and GPT-3.5 Turbo) do not reach consensus in any of the simulations and only reach partial levels of agreement with $|m|<0.5$. We show in Fig. 1 the evolution of the consensus level in societies of $N=50$ LLM agents and five different models, where the boxplot shows $C(t=10)$ over $20$ realizations. The two most advanced models considered, Claude 3 Opus and GPT-4 Turbo, reach consensus in all simulations, which corresponds to $|m|=1$ in the boxplot. On the other hand, less advanced models (Claude 3 Haiku and GPT-3.5 Turbo) do not reach consensus in any of the simulations and only reach partial levels of agreement with $|m|<0.5$. Finally, Llama 3 70B, a model with intermediate capabilities, displays mixed behavior, reaching consensus in only a fraction of the simulations.
We can get a deeper understanding of the underlying opinion dynamics by looking at the adoption probability $P(m)$, defined as the probability of an agent to support the norm as function of the average group opinion $m$. The right panel of Fig. 1 shows $P(m)$ for ten of the most popular LLMs and $N=50$ agents. The adoption probability is an increasing function of $m$ that approaches $1$ when $m=1$ and zero when $m=-1$. The most advanced models (GPT-4 family, Llama 3 70B, Claude 3 Sonnet and Opus) show a stronger tendency to follow the majority, with more pronounced S-shaped curves. On the other hand, less advanced models (GPT-3.5 Turbo, Claude 3 Haiku), have a weaker tendency to follow the majority, with GPT-3.5 Turbo going to some extent against the majority for small values of $m$.
@@
$P(m)=\frac{1}{2}[\tanh(\beta\cdot m)+1].$ (1)
@@
The parameter $\beta$, the majority force, gauges the inverse level of randomness in agent’s choices. For $\beta=0$ then $P(m)=1/2$: each agent behaves fully randomly (the new opinion is selected by coin-tossing) and consensus can be reached only by chance. This means that the expected time to reach consensus grows exponentially with $N$. For $\beta=\infty$ agents always align with the global majority and consensus is reached very quickly, on times growing logarithmically with $N$ [41].
The behavior of LLM agents is described well by (1), where fitting the parameter $\beta$ leads to high agreement with agent behavior for all agents except GPT-3.5 Turbo, for which no majority force is observable. This can be observed on $P(\tilde{m})$ as a function of the rescaled average opinion $\tilde{m}=m\cdot\beta$, shown on the right inset of Fig. 1, where all adoption probabilities (except GPT-3.5 Turbo) collapse on the same curve. Remarkably, the adoption probability of (1) is analogous to the behavior of a ferromagnet in the Curie-Weiss (CW) model evolving according to the Glauber dynamics [42, 43], akin to the probability for a spin to be up when the magnetization is $m$. In the CW model, there is a transition point for $\beta_{c}=1$ where order emerges (continuously) for $\beta>\beta_{c}$ [43], which implies that consensus will be feasible among LLM agents when the majority force is sufficiently above $1$.
Factors of the majority force
Figure 2: Role of Size and Language Understanding. Left: Plot of the majority force $\beta$, determining the tendency of LLM agents to conform to the majority versus the MMLU benchmark, measuring the language understanding capabilities of LLMs. The two quantities show a correlation of $0.76$, with the most capable models characterized by a stronger majority force. Right: $\beta$ as a function of the group size $N$ for various models. The majority force decreases in larger groups of LLM agents. The horizontal dashed line corresponds to $\beta=1$, the transition point of the Curie-Weiss model below which consensus is unfeasible.
The fit of all adoption probabilities to the function $P(\tilde{m})$ highlights that the difference in consensus formation between models is captured by the majority force parameter $\beta$. The left panel of Fig. 2 shows the values of $\beta$ for $N=50$ versus the score of each model in the MMLU benchmark, which measures the language understanding and cognitive capabilities of LLMs [44]. There is a clear growing relationship of $\beta$ with MMLU score, with a correlation coefficient of $0.76$ (p $<0.05$). This means that models with higher language understanding capabilities tend to exhibit a stronger tendency towards consensus, but none of the models shows consistent behavior against the majority. One might suspect that the majority force is actually dependent on the context window length as a plain “memory size”, but $\beta$ has a weaker, non-significant correlation with context window length ($0.49$, p $>0.1$, see SI for more details). Thus, language understanding is a stronger factor in the majority force of LLMs than context window length.
In animal (human and non-human) societies, group size plays a crucial role, with a progressive loss of norm stability as the number of individuals increases [30]. We hypothesize that a similar phenomenon may occur also in LLM societies, where agents might have their ability to coordinate limited by the size of their group. The right panel of Fig. 2 shows the estimated $\beta$ as a function of $N$ for various LLMs.  [^4] Given the cost of performing simulations with proprietary models, we analyzed increasing values from $N=10$ and stopped the analysis when $\beta$ reached values clearly below $1$, also excluding Claude 3 Opus in this analysis of $N$ due to its high cost. Independently of the model, there is a clear tendency for sufficiently large $N$: the larger the group, the weaker the majority force $\beta$. Models with a low MMLU tend to reach values of $\beta$ close to 1 for low $N$ of the order of few tens of agents, while some of the most advanced models still present a substantial majority force above $1$ even for $N=1000$.
Critical Consensus Size
@@
Figure 3: Critical Consensus Size. Left: Average time to reach consensus as a function of the group size for Llama 3 70B, GPT-4 Turbo, GPT-4o and the Curie-Weiss (CW) model with $\beta=3.75$. While both the CW model and GPT-4 Turbo display a slow growth of the consensus time, both Llama 3 70B and GPT-4o exhibit a rapid growth when the group size gets close to the critical consensus size. Averages are computed over $5$ realizations ($3$ for $N=500$ and only $1$ for $N=1000$ in the case of GPT-4 Turbo) and error bars show the range of values. Right: Critical consensus size $N_{c}$ of LLMs after which consensus gets (exponentially) unlikely. For humans, we report Dunbar’s number, while for GPT-4 Turbo and Claude 3.5 Sonnet we can only report a lower bound, since for $N$ up to $1000$ the majority force is well above the critical value $\beta_{c}=1$. The solid line represents an exponential fit of all points, excluding GPT-4 Turbo, Claude 3.5 Sonnet and humans. It shows that the latter models deviate from such a trend, exhibiting an $N_{c}$ value much larger than the scale of human group consensus.
As mentioned above, the adoption probability curves of the LLMs are the same as those for the CW model, where consensus is feasible only for $\beta>1$. However, this critical value $\beta_{c}=1$ is valid in the limit of very large groups, while here we are dealing with relatively small sizes compared to physical systems. Nevertheless, this transition is observable in our simulations of LLM societies when inspecting the mean time to reach consensus from a random state as a function of $N$. If $\beta$ were infinite, the consensus time would grow logarithmically with $N$ [41]. Instead, as shown in Fig. 3 using Llama 3 70B and GPT-4o as examples, the consensus time shows two different regimes. When $N$ is small the consensus time increases slowly, while for larger $N$, consensus time grows extremely fast. The transition between the two regimes occurs for a critical size $N_{c}$.
We can estimate this value from the size dependence of $\beta$ by setting $\beta(N_{c})=\beta_{c}\approx 1$. For Llama 3 70B $N_{c}\approx 50$, while for GPT-4o $N_{c}\approx 150$ (see Fig. 2). Note that, as detailed in Methods, this value of $N_{c}$ is actually an upper bound of the true value and the transition is expected as soon as $\beta$ gets close to one. On the other hand, a model with higher majority force like GPT-4 Turbo does not display any deviation from the scaling of the Curie-Weiss model for group size up to $N=1000$ (Fig. 3), showing a close adherence with the theory.
@@
The analogy with humans leads to the expectation that $N_{c}$ depends on the language understanding capabilities of LLMs, just as primates exhibit a growing relationship between neocortex ratio and average group size [30]. By studying the values of $\beta$ as a function of $N$ we can determine, for each model, the critical consensus size $N_{c}$, where $\beta\approx 1$. This is the size above which a group of LLMs does not reach consensus spontaneously over reasonable times. Examples of these values are plotted on the right panel of Fig. 2. In the case of GPT-4 Turbo and Claude 3.5 Sonnet, this provides a lower bound for this quantity as our analysis could not find values $\beta\approx 1$, even for large groups with $N=1000$. Note that we did not report GPT-4 since it would completely overlap with GPT-4 Turbo. We also plot, in the same figure, the critical consensus size for humans as Dunbar’s limit $N_{c}\approx 200$ [30]. LLMs display a similar trend as observed in anthropology, with language understanding capability predicting the limit of consensus size, i.e. the size of groups above which consensus becomes unlikely. The simplest LLMs show good agreement with exponentially growing $N_{c}$ with values for humans very close. This suggests that the reaching of consensus in groups is similarly related, both in humans and in LLMs, to cognitive capabilities, as measured by language understanding ability. On the other hand, some of the most modern and advanced LLMs go beyond this exponential scaling, reaching a super-human coordination capability despite having a human-level MMLU performance. For instance both GPT-4, GPT-4 Turbo and Claude 3.5 Sonnet are well above $\beta_{c}$ also for $N=1000$, the largest system we considered and substantially beyond Dunbar’s number.

### v4

@@ Abstract
Large language models (LLMs) are increasingly deployed in collaborative tasks involving multiple agents, forming an ”AI agent society" where agents interact and influence one another. Whether such groups can spontaneously coordinate on arbitrary decisions without external influence—a hallmark of self-organized regulation in human societies—remains an open question. Here we investigate the stability of groups formed by AI agents by applying methods from complexity science and principles from behavioral sciences. We find that LLMs can spontaneously form cohesive groups, and that their opinion dynamics is governed by a majority force coefficient, which determines whether coordination is achievable. This majority force diminishes as group size increases, leading to a critical group size beyond which coordination becomes practically unattainable and stability is lost. Notably, this critical group size grows exponentially with the language capabilities of the models, and for the most advanced LLMs, it exceeds the typical size of informal human groups. Our findings highlight intrinsic limitations in the self-organization of AI agent societies and have implications for the design of collaborative AI systems where coordination is desired or could represent a treat.
@@
Large Language Models (LLMs) have proven individual capabilities for a wide range of applications, such as summarization [13], sentiment analysis [32, 3], scientific research [7] or mathematical reasoning [37]. Agents driven by LLMs can be used in group settings where several agents interact with each other in collaborative tasks [26, 30, 39]. Collaboration setups where multiple LLMs have different roles and tasks, such as AutoGPT  [^1] https://github.com/Significant-Gravitas/AutoGPT, Microsoft’s AutoGen [42] or OpenAI SWARM [^2] https://github.com/openai/swarm, are qualitatively different from ensembling techniques, where interaction between different models is absent [27]. The interest for Agentic AI is on the rise, as also testified by the recent release of the Agent2Agent (A2A) protocol by Google  [^3] https://github.com/google/A2A, with the aim of simplify the integration of several AI agents from different frameworks. Recent advancement in on-device LLMs is also leading to AI-powered devices and assistants, such as Siri [^4] https://openai.com/index/openai-and-apple-announce-partnership/ or the Humane AI Pin, that can perform everyday tasks in interaction with each other, for example coordinating events or negotiating prices.
@@
As we move towards a society where AI agents interact with each other on our behalf, it becomes important to understand their ability to coordinate decisions in large groups. This can motivate new applications, but also help identify risks stemming from undesired collective behavior. For example, trading bots interacting through the stock market can lead to flash crashes [28] that motivate regulation like trading curbs. Current research on the behavior of LLMs has mostly focused on their behavior in isolation [1, 2, 17, 6, 36, 40] and collective behavior has been explored less [25, 4, 31], mostly with a focus on social simulation of network structures [16, 33, 12], opinion and information spreading [34, 14, 11] and online interaction [41, 35, 38]. To assess the reliability and trustworthiness of large numbers of interacting AI agents, we need to understand if they can form stable groups, display emerging coordination, what determines the abilities of AI agents to coordinate, and at what scale this can happen.
@@
Group coherence is related to coordination, as the agreement on shared decisions or norms creates a social contract that regulates behavior [20]. Taking quick collective decisions is often a necessity, even in situations where there is no information about the quality or utility of any option. This is the case, for example, of animal coordination when moving collectively and escaping predators [15]. Across species, this leads to a scaling of the average group size with brain structure, with human groups reaching sizes between 150 and 300 [18, 19], as documented by archaeological records [8] and contemporary experience [44]. To reach larger scales, human societies have built institutions and other ways of decision making, but the cognitive limit of about 250 contacts remains even in an online society [24, 21]. Such results lead to the hypothesis that intelligence and cognitive capabilities are a factor in the ability of AI agent societies to form stable cohesive groups. This application of concepts and methods from the analysis of human societies to the analysis of AI agent societies is a kind of AI anthropology [^5] In this piece, we adopt language that has a certain degree of anthropomorphising as an abstraction to explain AI agents as in https://huggingface.co/blog/ethics-soc-7. We do not mean that AI has a mind of itself but find this language useful to explain complex topics. that can leverage the insights of one discipline in another.
Figure 1: Emergence of stable groups. Simulation of a 150 AI agents powered by Llama 3 70B undergoing iterative opinion formation and splitting. The tree diagram (center) shows how the initial unstable large group progressively fragments into smaller stable cohesive groups through 10 iterations. We consider a group as stable if it does not split after two iterations. Circle sizes represent group population, with numbers indicating agent count. The left graph shows the average opinion over time for the initial 150-agent group. The right graph displays the average opinion for the 25-agent subgroup, which quickly fully coordinates.
@@
In this article, we investigate the ability of AI agents to form stable groups, coordinating about decisions for which there is no information supporting one option over another. The emergence of coordination is a foundational aspect of social systems, where individual interactions lead to the formation of a unified agreement or shared understanding without the need for a central authority or structure [22, 5]. We develop a framework to test if AI agent societies can form stable groups and up to which scale, and use it to analyze a benchmark of proprietary and open-source models. By applying insights from statistical physics, we measure a majority force that determines the possibility of coordination in societies of AI agents. This majority force is a function of cognitive capabilities of models and of group size, with coordination not emerging beyond a critical size for a given LLM. Furthermore, we find that AI agents running some of the most capable LLMs are able to coordinate at scales beyond what human groups typically achieve.
@@ Results
Stability of groups of AI agents
To investigate the coordination abilities of AI agent societies, we perform simulations using agents guided by various LLMs, belonging to the GPT, Claude, and Llama families. Simulations run as follows (see Methods for more details). We start with an initial group with $N$ agents. Each agent is assigned an initial opinion randomly chosen from a binary set (e.g., "Opinion A" and "Opinion B"). We then simulate a decision process with binary options, where there is no reason to favor the former over the latter, and vice versa. At each time step, a single agent is randomly selected to update their opinion. The selected agent receives the list of all other agents with their current opinions and is then prompted to choose their new opinion based on this information. This approach mirrors binary opinion dynamics such as the voter model or Glauber dynamics [10], where agents update their opinions based only on peer interactions. However, unlike traditional Agent-Based Models with predefined opinion update rules and equations, here AI agents autonomously decide their opinions based on their LLM. Agents are given some time too coordinate (in our simulations $t=10$, meaning that each agent is updated, on average, $10$ times) and then the simulation is stopped and the state of the system is evaluated. If agents coordinate, i.e., they all agree on the same opinion, the group remains cohesive. On the other hand, if both opinions survive and the AI agents do not all take the same decision, the group splits according to the opinion of the agents. The process is then iterated for the new groups that formed. We stop the evolution of the groups if they remain cohesive for $2$ iterations in a row, and stop the simulation when all groups have stopped this way. This process mimics a group of individuals that undergoes a series of important decisions, each potentially leading to a split. We show in Fig. 1 the result of such a process for Llama 3 70B and initial group size of $N=150$. As can be seen, the group is initially unstable and splitting step after step. However, the groups that result from these splits, when sufficiently small, show stability and do not split further. This shows that while large groups of AI agents are unstable, on smaller scales LLMs are able to coordinate and form cohesive social groups.
To characterize the evolution of the system, we need to focus on the dynamics within single groups. To this extent, we define the average group opinion
@@
Here $s_{i}$ is the opinion of agent $i$, the first opinion is $s_{i}=+1$ and the second is $s_{i}=-1$ (i.e., in favor and against the first option). $N_{+}$ and $N_{-}$ are the number of agents supporting the first and second opinion respectively, while $N$ is the total number of agents. In these terms we can define the coordination level $C=|m|$ that quantifies the level of agreement among agents. Full coordination corresponds to $C=1$, while $C=0$ means that the system is split in two groups of equal size and opposite opinion, thus coordination in absent. Partial coordination is achieved for intermediate values.
Figure 2: Coordination simulations among AI agents. Evolution of the coordination level over time for five different models and group size $N=50$. The box plots on the right show the final coordination level over $20$ simulations. Some models always fully coordinate, while the others never do so.
We show in Fig. 2 the evolution of the coordination level in societies of $N=50$ agents and five different models, where the boxplot shows $C(t=10)$ over $20$ realizations. The two most advanced models considered, Claude 3 Opus and GPT-4 Turbo, coordinate in all simulations, which corresponds to $|m|=1$ in the boxplot. On the other hand, less advanced models (Claude 3 Haiku and GPT-3.5 Turbo) do not reach coordination in any of the simulations and show an erratic behavior in their coordination level. Finally, Llama 3 70B, a model with intermediate capabilities, shows a clear tendency toward full coordination but does not reach within 10 iterations.
The majority force and its determinants
We can get a deeper understanding of the underlying opinion dynamics by looking at the adoption probability $P(m)$, defined as the probability of an agent to be in favor of the first option as function of the average group opinion $m$. The left panel of Fig. 3 shows $P(m)$ for ten of the most popular LLMs and $N=50$ agents. The adoption probability is an increasing function of $m$ that approaches $1$ when $m=1$ and zero when $m=-1$. The most advanced models (GPT-4 family, Llama 3 70B, Claude 3 Sonnet and Opus) show a stronger tendency to follow the majority, with more pronounced S-shaped curves. On the other hand, less advanced models (GPT-3.5 Turbo, Claude 3 Haiku), have a weaker tendency to follow the majority, with GPT-3.5 Turbo going to some extent against the majority for small values of $m$.
@@
$P(m)=\frac{1}{2}[\tanh(\beta m)+1].$ (1)
@@
This is a good fit in all cases except GPT-3.5 Turbo. This can also be seen by looking at the collapse plot showing $P(\tilde{m})$ as a function of the rescaled average opinion $\tilde{m}=m\beta$, shown in the inset of Fig. 3. All adoption probabilities (except GPT-3.5 Turbo) collapse on the same curve showing a universal behavior. The parameter $\beta$, the majority force, gauges the tendency to follow the majority versus randomness in agent’s choices. For $\beta=0$ then $P(m)=1/2$: each agent behaves fully randomly (the new opinion is selected by coin-tossing) and coordination can be reached only by chance. This means that the expected time to coordinate grows exponentially with $N$. For $\beta=\infty$ agents always align with the global majority and coordination is achieved very quickly, on a timescale growing logarithmically with $N$ [9].
Remarkably, the adoption probability in (1) is equal to the probability for a spin to be up when the magnetization is $m$ in the Glauber dynamics [23, 29] for the Curie-Weiss (CW) model, mean-field version of Ising model. The binary opinions can indeed be mapped to ±$\pm 1$ spin variables and the majority force to an inverse temperature. The equilibrium collective opinion $m^{*}$ can then be obtained from the self-consistency equation
$m^{*}=\tanh(\beta m^{*})$
from which the existence of a transition value $\beta_{c}=1$ can be derived (see Methods). This is a second order phase transition, where order emerges (continuously) for $\beta>\beta_{c}$ [29].
Wether or not a group of AI agents can coordinate will then depend on the value of $\beta$, on the corresponding equilibrium value of the collective opinion $m^{*}$ and on the fluctuations around this value. As we detail in the Methods section, we can identify three different regimes:
  • Absence of Coordination The majority force $\beta$ is below the critical value $\beta_{c}=1$. The equilibrium collective opinion is null, so the coordination level $C$ keeps fluctuating close to zero. This means that order is completely absent and the group is constantly in an uncoordinated state;
  • Partial Coordination The majority force satisfies $\beta_{c}<\beta<\beta_{t}$, where $\beta_{t}\approx(\log N)/2$. The equilibrium collective opinion $m^{*}$ is larger than zero, but smaller than one. Statistical fluctuations are not strong enough for the system to coordinate. As a consequence $C$ fluctuates around a value greater than 00 without ever reaching $1$;
  • Coordination For $\beta>\beta_{t}$ the equilibrium collective opinion is close to one and fluctuations are enough to coordinate. The coordination level quickly converges to $C=1$.
The different behaviors observed in Fig. 2 can then be explained as a direct result of the various LLMs having different majority forces, some of which below the critical value $\beta_{c}$ or the threshold value $\beta_{t}$. More details are reported in Methods.
Figure 3: Adoption probability and majority force. Left: adoption probability $P(m)$ as function of the collective opinion $m$ in a group of size $N=50$. Solid lines are fits of the curve $P(m)=0.5[\tanh(\beta\cdot m)+1]$ to the empirical data. The inset shows rescaled probabilities $P(\tilde{m})$ ($\tilde{m}=\beta m$) and confirms that all LLMs follow the same universal function. Right: majority force $\beta$ as a function of the group size $N$ for various models. The majority force decreases in larger groups of AI agents. The horizontal dashed line corresponds to $\beta_{c}=1$, the transition point of the Curie-Weiss model below which no order is present. The growing dot-dashed line indicates $\beta_{t}(N)$, the crossover point above which fluctuations lead to coordination.
The majority force does not depend only on the type of LLM. The right panel of Fig. 3 shows the estimated $\beta$ as a function of $N$ for various LLMs. We also show the three different regions in which the $\beta-N$ plane is divided and that corresponds to the three different regimes we discussed above.  [^6] Given the cost of performing simulations with proprietary models, we analyzed increasing values from $N=10$ and stopped the analysis when $\beta$ reached values clearly below $1$, also excluding Claude 3 Opus in this analysis of $N$ due to its high cost. For most models, there is a tendency for sufficiently large $N$: the larger the group, the weaker the majority force $\beta$. It is important to remark that this behavior is connected to the social aspect of the simulation and the prompt provided to the LLMs. No major performance difference is indeed observed in simple counting tasks when varying model or length. Details are reported in the SI.
In animal (human and non-human) societies, the size of the group plays a crucial role, with a progressive loss of stability of informal groups as the number of individuals increases [18]. Fig. 3 suggests that a similar phenomenon may occur also in AI agent societies, with the ability of agents to coordinate limited by the size of their group. This would explain the behavior observed in Fig. 1, with stability emerging in small but not in large groups.
Critical Group Size
@@
Figure 4: Critical Group Size. Left: Average time to coordinate as a function of the group size for Llama 3 70B, GPT-4 Turbo, GPT-4o and the Curie-Weiss (CW) model with $\beta=3.75$. While both the CW model and GPT-4 Turbo display a slow growth of the coordination time, both Llama 3 70B and GPT-4o exhibit a rapid growth when the group size gets close to the critical group size. Averages are computed over $5$ realizations ($3$ for $N=500$ and only $1$ for $N=1000$ in the case of GPT-4 Turbo) and error bars show the range of values. Right: Critical group size $N_{c}$ of LLMs after which coordination gets (exponentially) unlikely. For humans, we report Dunbar’s number, while for Claude 3.5 Sonnet we can only report a lower bound, since for $N$ up to $1000$ the majority force is well above the threshold value $\beta_{t}(1000)\approx 3.5$. The solid line represents an exponential fit of all points, excluding Claude 3.5 Sonnet and humans. The figure shows that the most capable models exhibits an $N_{c}$ value much larger than the scale of human informal groups.
We can characterize more precisely the transition from stable to unstable groups by inspecting, as a function of $N$, the mean time to coordinate starting from random initial opinions. As shown in Methods, this time grows with $N$ differently, depending on whether $\beta$ is larger or smaller than $\beta_{t}$. If $\beta>\beta_{t}$, the coordination time grows logarithmically with $N$ [9], meaning that even a very large group is stable and requires a short time to coordinate. This is the case for GPT-4 Turbo, as shown in Fig. 4, which perfectly follows the theoretical prediction of the CW model.
When $\beta<\beta_{t}$ instead the coordination time grows exponentially with $N$, implying in practice that large groups remain incohesive for any reasonable time interval. The behavior for Llama 3 70B and GPT-4o exhibits both regimes: when $N$ is small, the coordination time grows slowly, while for larger $N$ it diverges extremely fast. This confirms the presence of a size-induced transition occurring for some critical size $N_{c}$. We can estimate this value from the size dependence of $\beta$ by setting $\beta(N_{c})=\beta_{t}(N_{c})$. For Llama 3 70B $N_{c}\approx 30$, while for GPT-4o $N_{c}\approx 80$ (see Fig. 3).
@@
The analogy with primates leads to the expectation that $N_{c}$ depends on the intelligence of the LLMs, just as primates exhibit a growing relationship between neocortex ratio and average group size [18]. By studying the values of $\beta$ as a function of $N$ we can determine, for each model, the critical group size $N_{c}$, where $\beta\approx\beta_{t}$. We can then compare these values with a measure of models’ intelligence. Here we focus on the score in the MMLU (Massive Multitask Language Understanding) benchmark. This is a comprehensive evaluation of a language model’s knowledge and reasoning abilities across 57 subjects, ranging from mathematics and law to history and medicine, using multiple-choice questions. Similar results are obtained using also other measures, as detailed in the SI. We show a comparison of MMLU vs $N_{c}$ on the right panel of Fig. 4. In the case of Claude 3.5 Sonnet, this provides a lower bound for this quantity as our analysis could not find values $\beta\approx\beta_{t}$, even for large groups with $N=1000$. We also plot, in the same figure, the critical group size for humans as Dunbar’s limit $N_{c}\approx 200$ [18]. LLMs display a similar trend as observed in anthropology, with intelligence predicting the limit of coordination size, i.e. the size of groups above which coordination becomes unlikely. AI agents show good agreement with an exponentially growing $N_{c}$ with humans also being very closely aligned to this trend. On the other hand, some of the most modern and advanced LLMs reach a super-human coordination capability despite having a human-level MMLU performance. This is the case for Claude 3.5 Sonnet, that is well above $\beta_{t}$ also for $N=1000$, the largest system we considered, and GPT-4 Turbo, for which $N_{c}\approx 1000$. These values are substantially beyond Dunbar’s number and suggest that the ability to coordinate in groups is an emergent property, similarly related, in both humans and in LLMs, to cognitive capabilities.